OpenStudy (anonymous):

Steps for 3(2x+11)=6x+33

OpenStudy (anonymous):

If you're trying to show they are equal, you need to multiply the 3 through on the left side. 6x+33=6x+33

OpenStudy (anonymous):

Simplifying 3(2x + 11) = 6x + 33 Reorder the terms: 3(11 + 2x) = 6x + 33 (11 * 3 + 2x * 3) = 6x + 33 (33 + 6x) = 6x + 33 Reorder the terms: 33 + 6x = 33 + 6x Add '-33' to each side of the equation. 33 + -33 + 6x = 33 + -33 + 6x Combine like terms: 33 + -33 = 0 0 + 6x = 33 + -33 + 6x 6x = 33 + -33 + 6x Combine like terms: 33 + -33 = 0 6x = 0 + 6x 6x = 6x Add '-6x' to each side of the equation. 6x + -6x = 6x + -6x Combine like terms: 6x + -6x = 0 0 = 6x + -6x Combine like terms: 6x + -6x = 0 0 = 0 Solving 0 = 0

OpenStudy (anonymous):

that help

OpenStudy (anonymous):

Yes thanks fellas((:

OpenStudy (anonymous):

is me right i did the whole work can u talk tomorrow

OpenStudy (anonymous):

Lol maybe . I just zoned outta class today an I'm like stupid rn

OpenStudy (anonymous):

u nee help another question

OpenStudy (anonymous):

8x-3=11-6x

OpenStudy (anonymous):

Simplifying 8x + -3 = 11 + -6x Reorder the terms: -3 + 8x = 11 + -6x Solving -3 + 8x = 11 + -6x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '6x' to each side of the equation. -3 + 8x + 6x = 11 + -6x + 6x Combine like terms: 8x + 6x = 14x -3 + 14x = 11 + -6x + 6x Combine like terms: -6x + 6x = 0 -3 + 14x = 11 + 0 -3 + 14x = 11 Add '3' to each side of the equation. -3 + 3 + 14x = 11 + 3 Combine like terms: -3 + 3 = 0 0 + 14x = 11 + 3 14x = 11 + 3 Combine like terms: 11 + 3 = 14 14x = 14 Divide each side by '14'. x = 1 Simplifying x = 1

OpenStudy (anonymous):

that help

OpenStudy (anonymous):

Yes but like can you make drawings on here??

OpenStudy (anonymous):

i need to go talk to u tomorrow bye messga me tomorrow